OpenStudy (anonymous):

Find the limit \[\lim_{y \rightarrow 2-} (y-1)(y-2)/y+1 \]

OpenStudy (anonymous):

do we use factorizing method here?

OpenStudy (turingtest):

just plug the number in, and think about which terms will be positive and negative just to the left of 2

OpenStudy (turingtest):

oh wait, what am I saying? just plug it in...

OpenStudy (anonymous):

1/3 .. right?

OpenStudy (anonymous):

but in the book the answer is 0 ?

OpenStudy (anonymous):

(2-1)(2-2)=0

OpenStudy (turingtest):

\[\lim_{y\to2^-}{(y-1)(y-2)\over y+1}\]look at the numerator to see why it's zero

OpenStudy (anonymous):

(1)(0)/(2+1) = 0/3 = 0

OpenStudy (anonymous):

oh yea right .. I forgot multiplication of numerator .. Thanks